OpenStudy (anonymous):

can anyone help me?.fresh grapes contain 80% water and dry grapes contain 10% water ,if the weight of the dry grapes is 250 kg.what was it total weight when it was fresh

OpenStudy (anonymous):

Your initial weight \(x\), is going to be \(1x = 0.8x+0.2x\)

OpenStudy (anonymous):

Your new weight y, is going to be \(y= 0.1x+0.2x = 0.3x\)

OpenStudy (anonymous):

250*10/100=25 kg fresh for dry grapes 250*80/100=45kg fresh for fresh grapes

OpenStudy (nincompoop):

???

OpenStudy (anonymous):

\[ 0.3x = 250 \]

OpenStudy (anonymous):

250*100/20=750 for fresh

OpenStudy (anonymous):

2500/3=833.33 wrong @wio

OpenStudy (anonymous):

i know the method but didnt understand clearly so i asked the another way

OpenStudy (anonymous):